# Can Contrastive Learning Improve Class-Imbalanced Diffusion Model?

### v1

@@ Abstract
Training data for class-conditional image synthesis often exhibit a long-tailed distribution with limited images for tail classes. Such an imbalance causes mode collapse and reduces the diversity of synthesized images for tail classes. For class-conditional diffusion models trained on imbalanced data, we aim to improve the diversity of tail class images without compromising the fidelity and diversity of head class images. We achieve this by introducing two deceptively simple but highly effective contrastive loss functions. Firstly, we employ an unsupervised InfoNCE loss utilizing negative samples to increase the distance/dissimilarity among synthetic images, particularly for tail classes. To further enhance the diversity of tail classes, our second loss is an MSE loss that contrasts class-conditional generation with unconditional generation at large timesteps. This second loss makes the denoising process insensitive to class conditions for the initial steps, which enriches tail classes through knowledge sharing from head classes. Conditional-unconditional alignment has been shown to enhance the performance of long-tailed GAN. We are the first to adapt such alignment to diffusion models. We successfully leveraged contrastive learning for class-imbalanced diffusion models. Our contrastive learning framework is easy to implement and outperforms standard DDPM and alternative methods for class-imbalanced diffusion models across various datasets, including CIFAR10/100-LT, PlacesLT, TinyImageNetLT, and ImageNetLT.
@@ 1 Introduction
Recent advances in diffusion models ddpm ; ddim have led to breakthrough in various generation tasks such as image generation zhang2023adding ; stable_diffusion ; ruiz2023dreambooth , video generation ho2022imagen ; ho2022video , image editing couairon2022diffedit ; kawar2023imagic , 3D generation poole2022dreamfusion , etc. These diffusion-based generative models rely on large-scale datasets for training, which often follow a long-tailed distribution with dominant data for head classes and limited data for tail classes. Similar to the class-imbalanced recognition model liu2022open and the class-imbalanced generative adversarial network tan2020fairgen ; rangwani2022improving ; rangwani2023noisytwins ; utlo , diffusion model also tends to overfit on tail classes leading to mode collapse. As shown in Fig. 1, the original diffusion model generates similar or even repetitive images for tail classes. We aim to increase the diversity of tail class images while maintaining the fidelity and diversity of head class images for diffusion models.
@@
We look into a highly related and crucial problem of long-tailed image recognition, for which contrastive learning is an effective method jiang2021self ; li2022targeted . We are curious whether contrastive learning can improve long-tailed diffusion models for image synthesis. Contrastive learning is a general learning paradigm of minimizing the distance of positive pairs while maximizing the distance of negative pairs, which has attracted lots of research interest for self-supervised representation learning chen2020simple . On the other hand, diffusion model has also emerged as an effective model for representation learning fuest2024diffusion ; baranchuk2021label . We explicitly incorporate contrastive learning regularization to the diffusion model to further enhance representation learning towards better image generation.
Figure 1: To inspect overfitting on tail classes, we visualize training images (leftmost column) and synthetic images (third to rightmost columns) that are top-10 nearst neighbors sorted by distances to the anchor training image. KNN is based on L2 distances of Inception V3 embeddings. For each example, the top row is the results of DDPM ddpm , and the bottom row shows our method. The nearest neighbors from DDPM show severe mode collapse to the training image, while our method generated more diverse images. More examples are shown in the appendix A.3.3.
While at first glance, integrating contrastive learning into diffusion models seems straightforward, identifying suitable positive and negative pairs and designing an effective loss function are far from obvious. Also, training class-conditioned diffusion models involves both conditional generation and unconditional generation at multiple timesteps, which is more complicated than a feed-forward CNN or transformer. We found the devil is in the details and designed effective contrastive losses tailored for long-tailed diffusion models.
@@
Our first loss is an InfoNCE loss infonce with negative samples only, serving as regularization for denoising diffusion probabilistic models (DDPM). Given that mode collapse during inference manifests as generated samples being overly similar, particularly for tail classes, we introduce unsupervised contrastive learning with negative samples to maximize the distances between generated images. Unlike supervised contrastive learning khosla2020supervised , unsupervised contrastive learning does not differentiate between embedding/representation for different classes. Instead, it distinguishes between representations of images of the same class, hence increasing image diversity.
@@
Our second loss is an MSE loss designed to contrast estimated noises from conditional and unconditional generation, which effectively minimizes the KL divergence between latent distributions for conditional and unconditional generation. While such conditional-unconditional alignment seems undesirable with less controllability, a critical aspect of our approach is that our MSE loss is weighted more for larger timesteps corresponding to the initial stage of the reverse process. Given the known overfitting in conditional generation for tail classes for both GAN and diffusion models, we are motivated by the success of conditional-unconditional alignment for long-tailed GAN utlo ; transitioncgan , which facilitates knowledge sharing between head classes and tail classes. Notably, unconditional GAN generation has been observed to achieve superior FID than conditional generation under limited data transitioncgan . We adapt conditional-unconditional alignment from GAN to diffusion models. Unlike the GAN-based method utlo aligns conditional generation and unconditional generation for low-resolution images exhibiting intra-class similarity, we leverage observed image similarity during the initial denoising steps for tail classes and head classes si2024freeu . Hence, we propose to match conditional generation with unconditional generation for large timesteps.
@@
While contrastive learning has been proposed for diffusion models to improve adversarial robustness ouyang2023improving as guidance during inference and to find semantically meaningful directions dalva2024noiseclr , we effectively leveraged contrastive learning as regularization losses for class-imbalanced diffusion models. Our work has positive societal impact for content generation of underrepresented groups.
@@
  • We propose Contrastive Learning for Diffusion Model (CLDM) with imbalanced data. Our proposed contrastive losses are easy to implement with a standard DDPM training pipeline.
  • Our first loss, InfoNCE, employs unsupervised contrastive learning loss with negative samples only, enhancing within-class diversity.
  • Our second novel MSE loss aligns unconditional generation and conditional generation for the initial steps in the denoising process, facilitating knowledge sharing between head and tail classes.
  • We improve the diversity of tail class images for conditional generation while maintaining the fidelity and diversity of head class images for multiple datasets and various resolutions, including CIFAR10LT, CIFAR100LT, PlacesLT, TinyImageNetLT, and ImageNetLT. We also improve results for unconditional generation.
@@ 2 Related Work
Class-imbalanced Image Generation Generative models such as VAE vae , GAN stylegan ; biggan , and diffusion models ddpm ; stable_diffusion generate inferior images for tail classes when trained with real-world data with a long-tailed distribution. It has attracted a lot of research interest ai2023generative ; tan2020fairgen ; utlo ; cbdm ; oclt to address this issue for various types of models. Many methods address the problem of class imbalance by augmenting training data for the tail classes. A VAE is fine-tuned on tail classes under a majority-based prior ai2023generative . It is observed that GAN tan2020fairgen can amplify biases, leading to tail classes to be barely generated during inference, highlighting that the fairness in GAN needs improvement. Recent work addresses class-imbalanced diffusion models cbdm ; diffrop ; oclt by regularization losses to align or separate the distributions of synthetic images and their corresponding latent representations across different classes. For example, CBDM cbdm loss minimizes the distance of estimated noise items between these two models, the original DDPM model and a second model trained with pseudo labels which form a uniform distribution. DiffROP diffrop attempts to combine contrastive learning with diffusion model by maximizing the distance of distributions between classes. We aim to reduce overfitting for tail classes, while DiffROP diffrop reduces overlapping between different classes. The first part of our loss, the InfoNCE loss with negative pairs only, is defined for any pair of images regardless of their classes. However, DiffROP only considers pairs of images of different classes as negative pairs without regularizing images of the same class. Conceptually, we generalize DiffROP by including extra negative pairs of images within each class, hence promoting intra-class diversity.
We are inspired by UTLO utlo for long-tailed generation with GAN, which shares the latent representations of conditional GAN with unconditional GAN and implicitly shares knowledge between the head class and tail class. The motivation with UTLO utlo is that low-resolution images from GAN are similar for head classes and tail classes. We observe a similar phenomenon for denoised images in the initial steps of the denoising process, which motivates our contrastive loss. We adapt the conditional-unconditional alignment method from GAN transitioncgan ; utlo to diffusion models.
@@
Contrastive Learning for Representation Learning Both unsupervised contrastive learning chen2020simple ; moco and supervised contrastive learning khosla2020supervised ; li2022targeted are approaches used in representation learning to produce feature embeddings by maximizing the similarity between positive pairs and minimizing the similarity between negative pairs. Unsupervised contrastive learning does not require labels and typically augments the same data to form positive pairs (chen2020simple, ; moco, ). Supervised contrastive learning incorporates class labels and pulls together all samples from the same class while pushing apart samples from different classes. Contrastive learning is an effective method for self-supervised learning moco and long-tailed recognition jiang2021self ; li2022targeted . We effectively utilized contrastive learning for class-imbalanced diffusion models. Our work is different from the contrastive-guided diffusion process ouyang2023improving , which aims to improve adversarial robustness. Another difference is that we focus on contrastive learning during training, while contrastive-guided diffusion process ouyang2023improving is about guidance for inference. While the diffusion model originally for generation tasks becomes an emerging technique for representation learning fuest2024diffusion ; baranchuk2021label , we directly integrate contrastive learning regularization to diffusion models for better generation on imbalanced data. Notably, contrastive loss with negative samples exclusively focuses on separating dissimilar instances in the embedding space, which we leverage to address mode collapse, as detailed in Sec . 3.1. BYOL (BYOL, ) is a contrastive learning method with MSE loss for positive pairs only similar to our loss formulation in Sec. 3.2.
@@ 3 Method
Figure 2: Overview of the proposed framework of contrastive learning for diffusion model (CLDM). The encoder of the noise estimation network $e(\cdot)$ encodes an image with noise $x_{t}$ to a low-dimensional latent $h$, which is decoded to noise $\epsilon$ by the decoder $d(\cdot)$. We increase the distance of negative pairs of unconditional encoded latents for different samples by an InfoNCE loss $\mathcal{L}_{nceclr}$ with negative samples only. We form positive pairs from unconditional generation and conditional generation for the same sample $x_{t}^{i}$ and utilize an MSE loss $\mathcal{L}_{mseclr}$ to minimize their distances at initial time steps.
We propose Contrastive Learning for Diffusion Model (CLDM). Fig. 2 gives an overview of our framework. As is common in contrastive learning, our framework involes both negative and positive pairs, as outlined in Sec. 3.1 and Sec. 3.2, respectively. Sec. 3.3 summarizes our overall framework.
3.1 Negative Pairs and InfoNCE loss
@@
We observed the synthesized images of DDPM for a tail class concentrated around the limited training images in the latent space, leading to the issue of mode collapse, as shown in the Fig. 1, and the Fig. 5 in the appendix A.3.1 with visualizing the distribution of synthetic images for the tail class. By applying the InfoNCE loss with negative samples only, we intend to increase the distance of each synthesized image from other images in the latent space, hence increasing the diversity of synthetic images in particular for tail classes.
@@
We consider InfoNCE loss with negative samples only, where the negative samples are based on noisy images from a mini-batch. Specifically, an encoder $e(\cdot)$ encodes an image with noise $x_{t}$ to a low-dimensional latent $h$, which is decoded to noise $\epsilon$ by a decoder $d(\cdot)$ in a denoising network. Our InfoNCE loss is defined for the latents $h$ with each image in a mini-batch treated as an anchor. All other images are treated as negative samples. The anchor sample $x_{t}^{i}$ and negative samples $x_{t}^{j}$ are fed into the encoder $e(\cdot)$ to get the latents $h_{anc}^{i}=e(x_{t}^{i})$ and $h_{neg}^{j}=e(x_{t}^{j})$. Our InfoNCE contrastive loss $\mathcal{L}_{nceclr}$ is defined as follows:
$\mathcal{L}_{nceclr}=-\frac{1}{|B|}\sum_{i\in B}{\log\frac{\pi^{i}_{anc}}{\pi^{i}_{anc}+\sum_{j\in{B},j\neq i}{\pi^{j}_{neg}}}},\text{with }\pi^{i}_{anc}=\exp(\frac{h_{anc}^{i}\cdot h_{pos}^{i}}{\tau}),\pi^{j}_{neg}=\exp(\frac{h_{anc}^{i}\cdot h_{neg}^{j}}{\tau}),$ (1)
@@
where $B$ denotes a mini-batch, and $\tau$ is a temperature for softmax which we keep $0.1$ as the default setting. Note that it is common to augment an anchor image $x_{t}^{i}$ to be a positive sample in many unsupervised contrastive learning methods. Here, we don’t apply any data augmentation and directly treat the anchor latent $h_{anc}^{i}$ itself as the positive vector in the InfoNCE loss, i.e., $h_{pos}^{i}:=h_{anc}^{i}$.
3.2 Positive Pairs and MSE loss

### v2

# Contrastive Conditional–Unconditional Alignment for Long-tailed Diffusion Model

@@ Abstract
Training data for class-conditional image synthesis often exhibit a long-tailed distribution with limited images for tail classes. Such an imbalance causes mode collapse and reduces the diversity of synthesized images for tail classes. For class-conditional diffusion models trained on imbalanced data, we aim to improve the diversity and fidelity of tail class images without compromising the quality of head class images. We achieve this by introducing two simple but highly effective loss functions. Firstly, we employ an Unsupervised Contrastive Loss (UCL) utilizing negative samples to increase the distance/dissimilarity among synthetic images. Such regularization is coupled with a standard trick of batch resampling to further diversify tail-class images. Our second loss is an Alignment Loss (AL) that aligns class-conditional generation with unconditional generation at large timesteps. This second loss makes the denoising process insensitive to class conditions for the initial steps, which enriches tail classes through knowledge sharing from head classes. We successfully leverage contrastive learning and conditional-unconditional alignment for class-imbalanced diffusion models. Our framework is easy to implement as demonstrated on both U-Net based architecture and Diffusion Transformer. Our method outperforms vanilla denoising diffusion probabilistic models, score-based diffusion model, and alternative methods for class-imbalanced image generation across various datasets, in particular ImageNet-LT with $\times256$ resolution.
@@ 1 Introduction
Recent advances in diffusion models (Ho et al., 2020; Song et al., 2020) have led to breakthrough in various generation tasks such as image generation (Zhang et al., 2023; Rombach et al., 2022; Ruiz et al., 2023), video generation (Ho et al., 2022a; b), image editing (Couairon et al., 2022; Kawar et al., 2023), 3D generation (Poole et al., 2022), etc. These diffusion-based generative models rely on large-scale datasets for training, which often follow a long-tailed distribution with dominant data for head classes and limited data for tail classes. Similar to the class-imbalanced recognition model (Liu et al., 2022) and the class-imbalanced generative adversarial network (Tan et al., 2020; Rangwani et al., 2022; 2023; Khorram et al., 2024), diffusion model is not able to generate high-quality images for tail classes due to the scarcity of training data. As shown in Fig. 1, the original diffusion model with transformer backbone  (Ma et al., 2024) generates inferior images for a tail class (red wine) when trained using a long-tailed version of ImageNet. We aim to increase the fidelity and diversity of tail class images while maintaining the quality of head class images for diffusion models. Our work has a positive societal impact on content generation of underrepresented groups.
@@
We look into a highly related and crucial problem of long-tailed image recognition, for which contrastive learning is an effective method (Jiang et al., 2021; Li et al., 2022). Contrastive learning is a general learning paradigm of minimizing the distance of positive pairs while maximizing the distance of negative pairs, which has attracted lots of research interest for self-supervised representation learning (Chen et al., 2020). On the other hand, diffusion model has also emerged as an effective model for representation learning (Fuest et al., 2024; Baranchuk et al., 2021). We explicitly incorporate contrastive learning regularization to diffusion models to further enhance representation learning towards better image generation.
(a) SiT (Ma et al., 2024)
(b) SiT with CCUA (ours)
Figure 1: Synthetic images for a tail class (‘red wine’) by (a) standard SiT and (b) proposed CCUA framework. Both are trained on long-tailed ImageNet dataset for 160 epochs with 256x256 resolution.
@@
Our first loss is an unsupervised contrastive loss with negative samples only, serving as regularization for denoising diffusion probabilistic models (DDPM) and score-based diffusion models (SBDM). Given that mode collapse during inference manifests as generated samples being overly similar for tail classes, we introduce unsupervised contrastive learning with negative samples to maximize the distances between generated images. Unlike supervised contrastive learning (Khosla et al., 2020), unsupervised contrastive learning distinguishes between representations of images regardless of their class, hence increasing intra-class and inter-class image diversity.
@@
Our second loss is an alignment loss designed to align estimated noises from conditional and unconditional generation, which effectively minimizes the KL divergence between latent distributions for conditional and unconditional generation. While such conditional-unconditional alignment seems undesirable with less controllability, a critical aspect of our approach is that our alignment loss is weighted more for larger timesteps corresponding to the initial stage of the reverse process. Given the known overfitting in conditional generation for tail classes for both GAN, we are motivated by the success of conditional-unconditional alignment for long-tailed GAN (Khorram et al., 2024; Shahbazi et al., 2022), which facilitates knowledge sharing between head classes and tail classes. Notably, unconditional GAN generation has been observed to achieve superior FID than conditional generation under limited data (Shahbazi et al., 2022). We adapt conditional-unconditional alignment from GAN to diffusion models. Unlike the GAN-based method (Khorram et al., 2024) which aligns conditional generation and unconditional generation for low-resolution representations of images exhibiting intra-class similarity, we leverage observed image similarity during the initial denoising steps for tail classes and head classes (Si et al., 2024). Hence, we propose to match conditional generation with unconditional generation for large timesteps.
@@
While contrastive learning has been proposed for diffusion models to improve adversarial robustness (Ouyang et al., 2023), find semantically meaningful directions (Dalva & Yanardag, 2024), accelerate training (Yu et al., 2024), and regularize representation (Wang & He, 2025), we effectively leveraged contrastive learning for class-imbalanced diffusion models and demonstrate the superior performance of our method on long-tailed image generation via comprehensive experiments.
@@
  • We propose Contrastive Conditional-Unconditional Alignment for Diffusion Model (CCUA) with imbalanced data. Our proposed contrastive loss with conditional-unconditional alignment are easy to implement with standard DDPM and SBDM pipeline using both UNet-based architecture and Diffusion Transformer (DiT).
  • Our first loss, Unsupervised Contrastive Loss (UCL), employs unsupervised contrastive learning loss with negative samples only, enhancing within-class diversity.
  • Our second Alignment Loss (AL) aligns unconditional generation and conditional generation for the initial steps in the denoising process, facilitating knowledge sharing between head and tail classes.
  • We improved the diversity and fidelity of tail class images for conditional generation while maintaining the quality of head class images for multiple datasets and various resolutions, in particular ImageNet-LT with 256x256 resolution.
@@ 2 Related Work
Class-imbalanced Image Generation Generative models such as VAE (Kingma et al., 2019), GAN (Karras et al., 2019; Brock et al., 2019), and diffusion models (Ho et al., 2020; Rombach et al., 2022) generate inferior images for tail classes when trained with real-world data with a long-tailed distribution. It has attracted a lot of research interest (Ai et al., 2023; Tan et al., 2020; Khorram et al., 2024; Qin et al., 2023; Zhang et al., 2024) to address this issue for various types of models. Many methods address the problem of class imbalance by augmenting training data for the tail classes. A VAE is fine-tuned on tail classes under a majority-based prior (Ai et al., 2023). It is observed that GAN (Tan et al., 2020) can amplify biases, leading to tail classes to be barely generated during inference, highlighting that the fairness in GAN needs improvement. Khorram et al. (2024) propose a GAN-based long-tailed generation method, named UTLO, which shares the latent representations of conditional GAN with unconditional GAN and implicitly shares knowledge between the head class and tail class. The motivation with UTLO is that low-resolution representations of images from GAN are similar for head classes and tail classes. We observe a similar phenomenon for denoised images in the initial steps of the denoising process, and further propose a conditional-unconditional alignment loss designed for diffusion models. Recent work addresses class-imbalanced diffusion models (Qin et al., 2023; Yan et al., 2024; Zhang et al., 2024) by regularization losses to align or separate the distributions of synthetic images and their corresponding latent representations across different classes. For example, CBDM (Qin et al., 2023) loss minimizes the distance of estimated noise items between these two models, the original DDPM model and a second model trained with pseudo labels which form a uniform distribution. DiffROP (Yan et al., 2024) attempts to combine contrastive learning with diffusion model by maximizing the distance of distributions between classes. However, DiffROP only considers pairs of images of different classes as negative pairs without regularizing images of the same class. Conceptually, the first part of our method, the unsupervised contrastive loss designed to disperse unconditional latent space, generalizes DiffROP by including extra negative pairs of images within each class, hence promoting intra-class diversity.
@@
Contrastive Learning for Representation Learning Both unsupervised contrastive learning (Chen et al., 2020; He et al., 2020) and supervised contrastive learning (Khosla et al., 2020; Li et al., 2022) are approaches used in representation learning to produce feature embeddings by maximizing the similarity between positive pairs and minimizing the similarity between negative pairs. Unsupervised contrastive learning does not require labels and typically augments the same data to form positive pairs (Chen et al., 2020; He et al., 2020). Supervised contrastive learning incorporates class labels and pulls together all samples from the same class while pushing apart samples from different classes. Contrastive learning is an effective method for self-supervised learning (He et al., 2020) and long-tailed recognition (Jiang et al., 2021; Li et al., 2022). We effectively utilized contrastive learning for class-imbalanced diffusion models. Our work is different from the contrastive-guided diffusion process (Ouyang et al., 2023), which aims to improve adversarial robustness. Another difference is that we focus on contrastive learning during training, while contrastive-guided diffusion process (Ouyang et al., 2023) is about guidance for inference. While the diffusion model originally for generation tasks becomes an emerging technique for representation learning (Fuest et al., 2024; Baranchuk et al., 2021), we directly integrate contrastive learning regularization to diffusion models for better generation on imbalanced data. Notably, unsupervised contrastive loss with negative samples exclusively focuses on separating dissimilar instances in the embedding space, which we leverage to address mode collapse, as detailed in Sec . 3.1. In the context of generative modeling, REPA (Yu et al., 2024) aligns diffusion model features with features from a frozen vision encoder to accelerate training. Unlike REPA, we directly regularize diffusion model features without external models. Most relevant is concurrent work of dispersive loss (Wang & He, 2025), which is similar to our unsupervised contrastive loss with negative pairs only. Dispersive loss is directly and explicitly applied to conditional training, as designed for balanced diffusion models. By comparison, our unsupervised contrastive loss is formulated in the unconditional latent space and is implicitly extended to conditional training through our conditional–unconditional alignment loss. This carefully designed contrastive conditional-unconditional alignment framework for long-tailed diffusion models achieves better generation as shown in experiments.
@@ 3 Method
Figure 2: Overview of the proposed CCUA framework of contrastive learning for diffusion model. The noise estimation network is divided into two parts, a latent encode network e()e(*) and a decode network d()d(*). e()e(*) encodes an image with noise $x_{t}$ to a low-dimensional latent $h$, which is decoded to noise $\epsilon$ by d()d(*). Specifically, the latent encode network e()e(*) and decode network d()d(*) for unet-based model and diffusion transformer are shown in Fig. 5 in the appendix.A.1. We increase the distance of negative pairs of unconditional encoded latents for different samples by an unsupervised contrastive loss $\mathcal{L}_{ucl}$ with negative samples only, and align unconditional and conditional generation for the same sample $x_{t}^{i}$ and utilize an alignment loss $\mathcal{L}_{al}$ to minimize their distances at initial time steps.
We propose Contrastive Conditional-Unconditional Alignment (CCUA) framework for Diffusion Model. Fig. 2 gives an overview of our framework, while Fig. 5 in the appendix A.1 details how we apply CCUA framework into UNet-based model and Diffusion Transformer. Our framework involves a unsupervised contrastive loss with negative pairs only, and a conditional-unconditional alignment loss, as outlined in Sec. 3.1 and Sec. 3.2, respectively. Sec. 3.3 summarizes our overall framework.
3.1 Unsupervised Contrastive Loss
@@
We observed the synthesized images of DDPM for a tail class concentrated around the limited training images in the latent space, leading to the issue of mode collapse, as shown in the Fig. 9, Fig. 10 in the appendix A.3.3, and Fig. 8 in the appendix A.3.2 with visualizing the distribution of synthetic images for the tail class. By applying the contrastive loss with negative samples only, we intend to increase the distance of each synthesized image from other images in the latent space, hence increasing the diversity of synthetic images in particular for tail classes.
@@
We consider contrastive loss with negative samples only, where the negative samples are based on noisy images from a mini-batch. Specifically, we divide a noise estimation network into two parts, a latent encode network e()e(*), and decode network d()d(*). e()e(*) encodes an image with noise $x_{t}$ to a low-dimensional latent $h$, which is decoded to the estimate noise item $\epsilon$ by d()d(*). Our contrastive loss is defined for the latents $h$ with each image in a mini-batch treated as an anchor. All other images are treated as negative samples. The anchor sample $x_{t}^{i}$ and negative samples $x_{t}^{j}$ are fed into the encoder e()e(*) to get the latents $h_{anc}^{i}=e(x_{t}^{i};\emptyset)$ and $h_{neg}^{j}=e(x_{t}^{j};\emptyset)$. Our unsupervised contrastive loss $\mathcal{L}_{ucl}$ is defined as follows:
$\mathcal{L}_{ucl}=-\frac{1}{|B|}\sum_{i\in B}{\log\frac{\pi^{i}_{anc}}{\pi^{i}_{anc}+\sum_{j\in{B},j\neq i}{\pi^{j}_{neg}}}},\text{with }\pi^{i}_{anc}=\exp(\frac{h_{anc}^{i}\cdot h_{pos}^{i}}{\tau}),\pi^{j}_{neg}=\exp(\frac{h_{anc}^{i}\cdot h_{neg}^{j}}{\tau}),$ (1)
@@
where $B$ denotes a mini-batch, and $\tau$ is a temperature for softmax which we keep $0.1$ as the default setting. Note that it is common to augment an anchor image $x_{t}^{i}$ to be a positive sample in many unsupervised contrastive learning methods. Here, we don’t apply any data augmentation and directly treat the anchor latent $h_{anc}^{i}$ itself as the positive vector in the contrastive loss, i.e., $h_{pos}^{i}:=h_{anc}^{i}$.
Why contrastive loss encourages diversity and discourages mode collapse? In the worst case, that all embeddings collapse to a constant vector, the contrastive loss becomes $\log N$ for a batch with $N$ samples. Such constant embeddings yield the maximum possible loss and are therefore discouraged. Intuitively, contrastive loss introduces a repulsive force between different samples. The numerator remains constant, as self-similarity is always maximal, while the denominator aggregates pairwise similarities across the batch. Minimizing the loss requires reducing similarities between distinct samples, effectively pushing their embeddings apart in feature space. In this way, the model avoids collapse by maximizing inter-sample distances. Geometrically, the optimal solution (in the absence of augmentation) corresponds to embeddings being uniformly distributed on a hypersphere.
Batch Resample is a simple strategy for long-tailed recognition and generation (Zhang et al., 2024). However, it may lead to mode collapse due to repetitive images for tail classes. Since our unsupervised contrastive loss relies on pairs of images, batch resampling increases the chance of images from the same tail class appearing in the same batch, which further diversify tail class images. We choose batch resampling as an optional strategy for our framework and provide an ablation study in Sec. 4.3.
3.2 Conditional-unconditional Alignment Loss